You have $300,000 saved for retirement. Your account earns 9% interest. How much will you be able to pull out each month, if you want to be able to take withdrawals for 25 years?

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

Po=a(1-(1+(r/t)^(-nt)/(r/t)
r/k=.09/12=0.0075
Po=300000
therefore,
2250=a(1-(1+(r/t)^(-nt), nt=25*12=300
1-(1.0075)^-300=0.8937, don't round until the end
2250/.8937=a=$2517.59

Answer by ikleyn(52778)

.

           Let me disclose you a piece of info which can shock you.


9%  is very high annual percentage,  and having  $300,000  saved for retirement on your account,

it earns   0.09*300000 = 27,000  dollars annually.

So,  you can take  $27,000  from your account  (after it is compounded)  without decreasing this basic amount of  $300,000.


            AGAIN: under given condition,  you can have  $27000  once per year
            EVERY YEAR  without diminishing you base account  (!),  practically  FOREVER  (!)  (!)


Isn't it a shocking info ?


Consider this info as an entertainment.   I posted it to you to make you smiling . . .